OpenStudy (anonymous):

Fan and Medal Determine the equation of the line that passes through the points (2,-5) and (7,0). Use the form y=mx+b.

OpenStudy (arthur_ser):

The two points must evaluate the equation of the line, thus: -5=m*2 + b and 0=m*7 +b Solve the system by substracting each side: -5-0 =2m +b -7m -b -5=-5m m=1 Now put m into one of the previous equations to find b: 0=1*7+b b=-7 So the equation of the line is y=x - 7

OpenStudy (anonymous):

Thank you